40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 220 of 335  Not logged in ELOG logo
ID Date Author Type Category Subject
  5819   Sat Nov 5 01:10:29 2011 SureshUpdateIOOWFS output matrix measured (open loop)

 

The scripts and screens needed to make the MC WFS ouput matrix are once again functional

I corrected the WFS lockins' phases to ensure that the Q outputs are minimised.  Since all the lockins have the same relative phase with respect to the oscillator I found that the same phase works for all of them.  About 90 deg in this case.

The scripts used to make the WFS outmatrix measurement live in /cvs/cds/rtcds/caltech/c1/scripts/MC/WFS

1) setupWFSlockins:   This script makes sure that all the ASC, WFS_ LKIN and WFS_servo filter banks used in this measurement are set up properly.  It also sets the WFS_lockin oscillator to 10 Hz.  There are filter modules in the SIG filter bank of the WFS demodulators. 

2) senseWFSoutMATRX:  This script cycles through the various MC actuators ( MC 1 2 3 : PIT and YAW ) and measures the response of the various ASC sensors (WFS and MC2_TRANS QPD).

3) The data collected by the sensWFSoutMATRX can be analysed with a matlab file called " wfsmatrix3.m " located in a subdirectory under WFS called 'matlab'.   I have added some comments in this file to make it easier to follow.   The output of this file, at the moment, gives only the " Actuation Vectors " for WFS1P, WFS2P, WFS1Y and WFS2Y.  It ignores the MC2TransQPD for now. 

4)  The lockin outputs are given below ( the 'reduceddata' )

             wfs1p      wfs2p      mc2tp      wfs1y      wfs2y     mc2ty

mc1p    0.2926   -0.4086    0.2926    0.0340    0.0064    0.0001
mc2p   -0.2830   -1.3060   -0.2833    0.0628    0.1171   -0.0003
mc3p   -0.3283   -0.3455   -0.3288   -0.0456    0.0275    0.0000
mc1y    0.0440    0.0261    0.0429    0.7204    0.9351   -0.0008
mc2y   -0.1006    0.0850   -0.1036   -1.5509   -0.3882    0.0165
mc3y     0.0150   -0.0832    0.0144    0.1114   -1.0573    0.0006

5) The actuation vectors are given below

Pitch WFS1P WFS2P
MC1 1.00 -0.86
MC2 -0.12 -1
MC3 -0.72 0.09

 

Yaw WFS1Y WFS2Y
MC1 0.16 0.59
MC2 -1.00 0.20
MC3 0.51 -1

6) This measurement was performed with the WFS servo loops open. I will try to close the loops with this matrix and run the script again to measure the output matrix in closed loop.

7) This a.vectors obtained above are significantly different from that obtained a while ago (elog 5668) before the lockin demod phases (relative to each other) were fixed.  This could also be because both are open loop measurements and we might have wandered into the nonlinear regime of the WFS sensors.

 

 

 

  5818   Sat Nov 5 00:24:13 2011 SureshUpdateASCASS scripts gone

Quote:

Did somebody delete all the scripts in /opt/rtcds/caltech/c1/scripts/ASS ?

 I have moved all the MC_ASS scripts to a directory called MC under ASS

 

  5817   Sat Nov 5 00:04:23 2011 kiwamuUpdateASCASS scripts gone

Did somebody delete all the scripts in /opt/rtcds/caltech/c1/scripts/ASS ?

  5816   Fri Nov 4 21:52:58 2011 DenUpdateAdaptive Filteringcoherence

[Mirko, Den]

We still think about the coherence between seismic noise and mode cleaner length. We beleive that

1. Below ~0.1 Hz tilt affects on the seismometers as was simulated http://nodus.ligo.caltech.edu:8080/40m/5777

2. From 0.1 to 1 Hz is an interesting region. We try to figure out why we do not see any coherence here. Tilt does not seem to dominate.

At 1 Hz coherence might be lost because of the sharp resonance. For example, if the mirror is suspended to the platform by wires with f = 1 Hz and Q = 1000, then the coherence between platform motion and mirror motion will be lost as shown on the figure below.

mirror_platform.jpg

For this reason we tried to "help" to the adaptive filter to guess transfer function between the ground motion and mirror motion by multiplying seimometer signal by the platform -> mirror transfer function. As we do not know exactly eigen frequency and Q of the wires, we did a parametric simulation as shown on the figure below

coherence.jpg

The maximum coherence that we could achieve with treak was 0.074 compared to 0.056 without. This was achieved at f=1.0011 Hz but with surprisingly high Q = 330. And though this did not help, we should keep in mind the tecnique of "helping" the adaptive filter to guess the transfer function if we partly know it.

Another unexpected thing is that we see come coherence between gur1_x and mode cleaner WFS pitch signal at frequencies 0.1 - 1 Hz

Screenshot-4.png

 

 

 From this we can suggest that either mode MC_F channel does not completely reflect the mc length at low frequencies or WFS2 shows weard signal.

  5815   Fri Nov 4 21:15:38 2011 KatrinUpdateGreen LockingNew fiber coupler (YARM)

63% coupling efficiency into the new fiber collimator (Thorlabs XXXX) and the blue fiber.This should be sufficient for a beat measurement with the PSL laser.

I think the coupling efficiency is not too bad with having no mode matching lenses and no adjustable collimator lens.

 

252mW in front of the fiber

159 mW fiber output

  5814   Fri Nov 4 19:30:06 2011 ranaUpdateSUSDr. SUS paths updated--question of human oversight remains

My inclination is to not do the writing of the matrices automatically nor to do the weekly kicks. Its nice to have long locks of the MC, etc.

I suggest just making the kick on Sundays when someone intentionally asks for it (e.g. by pressing a button on Friday). The free-swinging ringdown ought to be a opt-in kind of feature, not opt-out.

  5813   Fri Nov 4 16:23:41 2011 SureshUpdateIOOMC spot decenter measured

After Kiwamu adjusted the MC2 PIT to accommodate the limited range of the PZT1 ( elog ),  I remeasured the spot positions today.  

  MC1P MC2P MC3P MC1Y MC2Y MC3Y
Yesterday 0.1354 -0.2522 -0.1383 -1.0893 0.7122 -1.5587
Today 4.0411 4.4994 3.5564 -1.4170 -0.2606 -1.7109

 

As expected there is a translation of the beam axis to one side (Up? Down?) .  

I wonder how a beam translation by 5mm solved the PZT1 angular range limitation problem (?!)

  5812   Fri Nov 4 15:26:54 2011 JenneUpdateLSCLSC model recompiled

I moved the place where we take the OAF Degree of Freedom signals from - now it's the error point rather than the feedback for DARM, CARM, MICH, PRCL, SRCL, XARM and YARM.  I didn't do anything to MCL.

While trying to compile, there was something wrong with the lockins that were there...it complained about the Q OUTs being unconnected.  I even reverted to the before-I-touched-it-today version of c1lsc from the SVN, and it had the same problem.  So, that means that whomever put those in the LSC model did so, and then didn't check to see if the model would compile.  Not so good.

Anyhow, I just terminated them, to make it happy.  If those are actually supposed to go somewhere, whoever is in charge of LSC lockins should take a look at it.

Also, as Mirko mentioned in the previous elog 5811, we wanted to calculate the effect on the MC without actuating, so we put in a new summing point and a filterbank so we have testpoints.

LSC model recompiled.

OAF model recompiled.

FB restarted because of the new channels added to OAF.

  5811   Fri Nov 4 15:24:13 2011 MirkoUpdateAdaptive FilteringAdaptive FF on the MC doesn't make sense

[Den, Jenne, Mirko]

DSC_3585.JPG

Here is the story:

1. High gain
The control loop has a high gain at the interesting frequencies. The error point (EP) before the servo is approx. zero and the information how much the mirror would move is in the feedback point (FB) behind the servo. The mirror doesn’t actually move because of the high gain. This is the case of the grav. wave detectors and medium frequencies (> approx. 50Hz, <<1kHz). Adding feed-forward (FF) to this doesn’t actually keep the mirror quieter. In fact if you look into the FB and subtract the seismic you make the mirror move more. Yes this is the case we have for the mode cleaner, doesn’t make sense.
In a real GW detector FF however isn’t totally useless. The FB tells you how much the mirror moves, due to GWs, seismic etc. When you record the FB and subtract (offline) the seismic you get closer to the real GW signal.

2. Low gain
When you, for technical reasons, can’t have a high gain in your control loop the EP contains information of how the mirror actually moves. You can then feed this into the adaptive filter and add its output to the FB. This will minimize the EP reducing the actual mirror motion. This is the case we will have for most or all other degrees of freedom in the 40m.

The reason we have so much gain in the mode cleaner length control is that we don’t actually move mirrors around. We change the frequency of the incoming laser light. You can do that crazy fast with a big amplitude. This gives us a high UGF and lots of gain in the 1Hz range we are interested in.

We now change the adaptive filter to look at the EP for all DOFs except for the MC. We calculate the effect of the FF on the MC length signal without ever applying the FF to the MC length control.

  5810   Fri Nov 4 14:18:24 2011 MIrkoUpdateAdaptive FilteringCoherence between seismometers and MC length

Looking into the coherence between the seismometers and IMC length (MC_F):

FIrst with the seismometers only AC filtered at around 0.003 Hz and AA30Hz:
Coherence_without_compensation.pngWith_Pend_compensation.mat

Ignore the increase in coherence at very low frequencies. That is an artefact.

Then with an additional filter single complex pole @1Hz Q=1000 (giving 20dB per decade in attenuation above 1Hz) , only for GUR1X:
Coherence_with_compensation.png

  5809   Fri Nov 4 13:53:33 2011 KatrinUpdateGreen LockingSHG temperature (YARM)

Quote:

I must confess that I changed the temperature of the laser via the dial yesterday. I believe the initial (displayed) temperature was ~19o, whereas it is now probably in the high 20s. Sorry.

Quote:

Changing the crystal temperature changes the laser frequency. This will causes the beat note missing at the vertex.

In other words, you will find the beat note at the vertex when the actual temperature of the crystal is reproduced as before,
no matter how the dial setting/temp voltage input is.

 

Okay, my elog entry was not clear I changed the temperature of the SHG which only changes the conversion efficiency. 

 

 

Anyhow, since the laser set temperature and thus the laser frequency has been changed by Zach and I couldn't find a note

of the original laser crystal temperature, my plan is to reset the SHG temperature to the old value, set the laser crystal temperature

around 19°C and do fine adjustment of that temperature by optimising the doubling efficiency. Okay?

 

  5808   Fri Nov 4 13:13:25 2011 ZachUpdateGreen LockingPDH box #17 modified, too

I also modified the Y-Arm PDH box itself slightly. Previously, there was a flying 10k resistor from the SWEEP input to TP2. I don't see the point of this, so I moved it from TP2 across R19 (to the same point where it is on the gyro PDH boxes) to allow for excitation signals to be injected with the loop closed (i.e., with the SWEEP switch off). This is useful for OLTF measurements.

  5807   Fri Nov 4 13:04:50 2011 ZachUpdateGreen LockingPassive summing box modifications

I spent some time the other day trying to diagnose the problem with the Y Arm universal PDH box (S/N 17), which Katrin has been unable to use for locking the green beam. As far as I can tell, there is nothing wrong with the box itself (though the weird TF behavior that Katrin noticed was not initially reproducible, so its cause may still be there).

I did notice that I was unable to generate a PDH error signal using the universal box. In this configuration, a summing circuit is needed to add the PZT modulation signal (fmod = 178875 Hz) in along with the feedback signal. To do this, Katrin was using a slightly different version of the passive summing box that Kiwamu built for the X Arm. I read this entry to understand how it is supposed to work and noticed that the "expected transfer functions" were not what the circuit actually does. I have talked to Kiwamu about it and he found that he posted the wrong TFs (he has the right ones on his computer). As you can see from the plot below, there is extra low passing that severely attenuates the modulation path to the PZT. In addition, there is a phase shift of ~-60 degrees, which is bad.

To combat this, I propose we simply change the resistor in the modulation path from 1M to 10k. This leaves the feedback path TF unchanged, and changes the mod path into a sort of bandpass filter for the modulation frequency. The fact that the phase is near zero at fmod means we don't have to come up with some way to phase shift the signal for demodulation. The attenuation level of ~-36 dB is also convenient: The ZAD-8 mixer wants 7 dBm, so, 10 dBm (FG) - 3 dB (splitter) - 36 dB (sumbox) = -25 dBm ~ 12 mV. This is roughly the desired PZT voltage level.

sum_circuit_new.png sumbox_TFs.png

  5806   Fri Nov 4 05:24:56 2011 kiwamuUpdateLSCoffset introcuded in MC and IFO-configure script modified

[Offsets in MC]

 I have introduce an offset in MC2 PIT because the PZT1 again started railing.

Right now the PZT1 EPICS value is within the range happily.

Please keep this MC eigen axis as a nominal configuration.

 

[IFO-configure script]

 I have modified the IFO configure scripts such that XARM and YARM are locked with POX11 and POY11 respectively.

A big advantage in use of POX and POY is that we don't need to misalign ITMs when we align each arm.

Those scripts are now available from the C1IFO_CONFIGURE screen as usual.

  5805   Fri Nov 4 00:25:49 2011 ZachUpdateSUSDr. SUS paths updated--question of human oversight remains
  • I have replaced all instances of /cvs/cds/rtcds with /opt/rtcds in the Dr. SUS code. 
  • We discussed placing a human in charge of whether or not the input matrices get written to the frontend. I am not sure if we reached a definitive conclusion. Currently, the code is set to write the matrices automatically. What to do?
    • If we decide that oversight is necessary, I suggest that we leave the publishing of the results to the elog intact. This way, it will be someone's job to read the weekly Dr. SUS diagnosis on Sunday night (or Monday morning), and run a simple script that writes the matrices. This seems like the most reasonable solution.

I await responses...

  5804   Fri Nov 4 00:13:24 2011 KojiUpdateLSCXARM locked with POX11

XARM lock was achieved by POX11_I

Summary:

- The  whitening gains of POX11_I and Q are 42dB so that POX11_I have the same amplitude as AS55_I

- The demod phase of POX11 was adjusted to eliminate the PDH signal from the Q phase. The phase is  -100.5deg.

- In order to lock the XARM with POX11_I_ERR, I had to increase the trigger threshold from 0.1 to 0.2 as the arm was
kicked with the threshold of 0.1.


Method

- Lock the X arm with AS55_I at the XARM configuration.

- Adjusted POX11 demod phase so that POX11_Q is minimized.

- POX I&Q whitening gains were adjusted. When they are 42dB, POX11_I_ERR and AS55_I_ERR have almost the same signal amplitude.
(In reality, POX11_I_ERR has +1dB larger amplitude.)

- Adjusted POX11 demod phase again with better precision.

- Measured transfer function between AS55_I_ERR and POX11_I_ERR. As the sign was opposite, the demod phase was -180deg flipped.

- Tried to lock the arm with POX11_I_ERR. It did not acquire the lock. The arm looked kicked by the servo.

- Increased the trigger threshold from 0.1 to 0.2. Now the arm is locked with POX11_I_ERR.

Attachment 1: POX11.pdf
POX11.pdf
  5803   Thu Nov 3 22:32:48 2011 SureshUpdateIOOMC realigned to center the spots on actuation nodes

Quote:

Several activities in the past week (elog1, Mirko's rough realignment of MC and adjustment of the PSL zig-zag on Monday and elog2 ), had led to an MC state where the spots were not centered on the actuation nodes.

The change in the C1IOO model part dealing with MC_ASS led to the disappearance of all MC_ASS filter definitions in the lockins.  I have fixed all that, remade the screens and scripts for making the MC Decenter measurements.

I have just completed the re-centering of the spots on the MC and this of course will lead to some change in the input pointing. 

The MC_REFL beam also was recented on the WFS and the MC2_TRANS QPD.

Will add the mc-decenter measurement in a few mins

 The current MC spot decentering is given below:

spot positions in mm:

MC1P     MC2P     MC3P      MC1Y     MC2Y     MC3Y

0.4089    0.4800   -0.1266   -1.4095    0.3808   -1.7517

 

The yaw measurement probably has the wrong scale factor in the conversion to mm.  It could be under estimated by a factor 2.65/2.00 since the 10% step in coil gains produces a 2mm offset rather than the expected 2.65mm.  See the figures below.  I will check this during the next iteration when another mode clearner alignment comes up.

As I had to redefine all the MC_ASS lockin filters it is possible that Lockin phase might have changed by a few degrees giving rise to a change in the scale factor.

20111103_1.png    20111103_2.png

  5802   Thu Nov 3 19:58:18 2011 kiwamuUpdateLSCPOX11 demod board : modification done

The modification on the POX11 demod board has been successfully done.

I followed the procedure which had been posted in a past entry (#4554).

The home-made splitter was replaced by PSCQ-2-51W, which has a relatively wide band of 5 - 50 MHz.

The usual orthogonality adjustment will be done in the daytime.

 

 The attached snapshot was taken when an sinusoidal RF signal with a slight frequency offset from LO was injected to the RF input.

It is clear that the I and Q output show healthy signals (i.e. almost the same amplitude and 90 deg phase difference.)

POX11demod.png

Quote from #5801

I am going to replace the splitter which had been made with a hand-wounded coil because it can work only at a specific tailored frequency.

  5801   Thu Nov 3 18:41:36 2011 kiwamuUpdateLSCPOX11 demod board : haven't been modified yet

I pulled out the POX11 demod board and found the power splitter on the board hadn't been modified yet.

I am going to replace the splitter which had been made with a hand-wounded coil because it can work only at a specific tailored frequency.

Quote from #5753
The POX11 demodulation board is broken

  5800   Thu Nov 3 17:24:59 2011 ZachUpdateGreen LockingSHG temperature (YARM)

I must confess that I changed the temperature of the laser via the dial yesterday. I believe the initial (displayed) temperature was ~19o, whereas it is now probably in the high 20s. Sorry.

Quote:

Changing the crystal temperature changes the laser frequency. This will causes the beat note missing at the vertex.

In other words, you will find the beat note at the vertex when the actual temperature of the crystal is reproduced as before,
no matter how the dial setting/temp voltage input is.

 

  5799   Thu Nov 3 17:19:54 2011 SureshUpdateIOOMC realigned to center the spots on actuation nodes

Several activities in the past week (elog1, Mirko's rough realignment of MC and adjustment of the PSL zig-zag on Monday and elog2 ), had led to an MC state where the spots were not centered on the actuation nodes.

The change in the C1IOO model part dealing with MC_ASS led to the disappearance of all MC_ASS filter definitions in the lockins.  I have fixed all that, remade the screens and scripts for making the MC Decenter measurements.

I have just completed the re-centering of the spots on the MC and this of course will lead to some change in the input pointing. 

The MC_REFL beam also was recented on the WFS and the MC2_TRANS QPD.

Will add the mc-decenter measurement in a few mins

  5798   Thu Nov 3 17:04:23 2011 KojiUpdateGreen LockingSHG temperature (YARM)

Changing the crystal temperature changes the laser frequency. This will causes the beat note missing at the vertex.

In other words, you will find the beat note at the vertex when the actual temperature of the crystal is reproduced as before,
no matter how the dial setting/temp voltage input is.

  5797   Thu Nov 3 16:43:44 2011 KatrinUpdateGreen LockingSHG temperature (YARM)

Plugging in the thermal feedback BNC cable to the laser reduced the DC voltage of the green PDH photo diode from 3.12 V to 1.5V off resonance.

The power emitted by the laser was the same as in the case without that cable. Note LT, i.e. measured crystal temperature, of the laser show a

different value when the BNC is connected, but the manual clearly states that this display does not work properly if a cable is connected to the

slow BNC plug, an offset is added.

The power of the 532nm light behind the SHG oven has been reduced from 1mW to 0.4mW. I changed the crystal temperature such that the power

of the green light is 1mW. With this new temperature setting the laser can be locked again.

 

  w/o BNC cable at slow plug w/ BNC cable at slow plug
T (°C) 29.776 29.776
LT (°C) 39.2 48.4
1064nm power (mW) 440 448
Temp. at TC200 (°C) 35.7 36.4
532nm power in front of Isolator (mW) 1.0 1.0

 

 

 

  5796   Thu Nov 3 16:31:47 2011 kiwamuUpdateLSCPDA10CF as POP22/110 : better

The POP22/110 RFPD has been replaced by PDA10CF. As a result the 22 and 110 MHz signals became detectable.

However the signal level maybe too low according to a quick look with an RF spectrum analyzer.

The level at 22 and 110 MHz were both approximately -70 dBm although these values were measured when the central part was freely swinging.

Perhaps we need to amplify the signals depending on the actual SNR.

 

 Also I have updated the optical tables' wiki page :

http://blue.ligo-wa.caltech.edu:8000/40m/Optical_Tables

Quote from #5794

I will replace it by PDA10CF, which is made from InGaAs and supposed to have 10 times bigger responsibility.

  5795   Thu Nov 3 15:14:22 2011 KojiUpdateCDSCSS/BOY installed on pianosa

How to run/use CSS/BOY

(PREPARATION)

0) Everything runs on pianosa for now.

1) type css to launch CSS IDE.

2) You may want to create your own project folder as generally everything happens below this folder.

=== How to make a new project ===
- Right-click on tree view of Navigator pane
- You are asked to select a wizard. Select "General -> Project". Click "Next".
- Type in an appropriate project name (like KOJI). Click "Finish"
- The actual location of the project is /home/controls/CSS-Workspaces/Default/KOJI/ in the above example

(PLAY WITH "Data Browser", THE STRIPTOOL ALTERNATIVE)

1) Select the menu "CSS -> Trends -> Data Browser". A new data browser window appears.

2) Right-click on the data browser window. Select the menu "Add PV". Type in the channel name (e.g. "C1:LSC-ASDC_OUT16")

3) Once the plot configuration is completed, it can be saved as a template. Select the menu "File -> Save" and put it in your project folder.

4) Everything else is relatively straightforward. You can add multiple channels. Log scaling is also available.
I still don't find how to split the vertical axis to make a stacked charts, but I don't surprise even if it is not available.
CSS_snapshot1.png

(HOW TO MAKE A NEW "BOY" SCREEN)

0) Simply to say BOY is the alternative of MEDM. The screen file of BOY is named as  "*.opi" similar to "*.adl" for MEDM.

1) To create a new opi file, right-click on the navigator tree and select the menu "New -> Other".

2) You are asked to select a wizard. Select "BOY -> OPI file" and click "Next".

3) Type in the name of the opi file. Also select the location of the file in the project tree. Click "Finish".

4) Now you are in OPI EDITOR. Place your widgets as you like.

5) To test the OPI screen, push the green round button at the top right. The short cut key is  "CTRL-G".
OPI_EDITOR.png

(HOW TO EDIT AN EXISTING OPI FILE)

1) Right click an OPI file in the navigator tree. Select the menu "Open WIth -> OPI Editor". That's it.

(HOW TO CONVERT AN EXISTING ADL FILE INTO OPI FILE)

1) You need to copy your ADL file into your project folder. In this example, it is /home/controls/CSS-Workspaces/Default/KOJI/

2) Once the ADL file is in the project folder, it should appear in the navigator tree. If not, right-click the navigator pane and select "Refresh".

3) Make sure "ADL Parser" button at the left top part is selected, although this has no essential function.
This button just changes the window layout and does nothing. But the ADL Tree View pane would be interesting to see.

3) If you select the ADL file by clicking it,the tree structure of the ADL file is automatically interpreted and appears in the ADL Tree View pane.
But it is just a display and does nothing.

4) Right-click the ADL file in the navigator pane. Now you can see the new menu "BOY". Select "BOY -> Convert ADL File to OPI".

5) Now you get the opi version of that file.
The conversion is not perfect as we can imagine. It works fine for the simple screens.
(e.g. matrix screens)
But the filter module screens get wierd. And the new LSC screen did not work properly (maybe too heavy?)
ADL2OPI1.png ADL2OPI2.png

(HOW TO RUN OUR SHELL/PERL/PYTHON SCRIPTS FROM BOY)

CSS has javascript/python scripting capability.
I suspect that we can make a wrapper to run external commands from python script although it is not obvious yet.

  5794   Thu Nov 3 14:25:52 2011 kiwamuUpdateLSCPDA10A as POP22/110 : too small signal

It turned out that the signal was too small with PDA10A to detect the 22 and 110 MHz RF sidebands.

The DC output coming out from it was about half mV or so (corresponding to few uW in laser power) when the PRCL was locked to the carrier.

This is because PDA10A is a silicon detector which is more sensitive to visible light than IR.

The reason we chose PDA10A was that it has relatively a large diode size of 1 mm in diameter.

However according to the data sheet the responsibility at 1064 nm is about 0.05 A/W which is sad.

I will replace it by PDA10CF, which is made from InGaAs and supposed to have 10 times bigger responsibility.

Though the diode size will be half mm in diameter, which may require another strong lens in front of it.

Quote from #5788
 The POP22/110 RFPD has been installed. It is PDA10A from Thorlabs instead of the usual home-made RFPD.
  + Fine alignment

  5793   Thu Nov 3 13:00:52 2011 JenneUpdatePhotosFormatting of MEDM screen names

Quote:

After lots of trial and error, and a little inspiration from Koji, I have written a new script that will run when you select "update snapshot" in the yellow ! button on any MEDM screen. 

Right now, it's only live for the OAF_OVERVIEW screen.  View snapshot and view prev snapshot also work. 

Next on the list is to make a script that will create the yellow buttons for each screen, so I don't have to type millions of things in by hand.

The script lives in:  /cvs/cds/rtcds/caltech/c1/scripts/MEDMsnapshots, and it's called....wait for it....... "updatesnap".

 Currently the update snapshot script looks at the 3 letters after "C1" to determine what folder to put the snapshots in.  (It can also handle the case when there is no C1, ex. OAF_OVERVIEW.adl still goes to the c1oaf folder).  If the 3 letters after C1 are SYS, then it puts the snapshot into /opt/rtcds/caltech/c1/medm/c1sys/snap/MEDM_SCREEN_NAME.adl

Mostly this is totally okay, but a few subsystems seem to have incongruous names.  For example, there are screens called "C1ALS...." in the c1gcv folder.  Is it okay if these snapshots go into a /c1als/snap folder, or do I need to figure out how to put them in the exact same folder they currently exist in?  Or, perhaps, why aren't they just in a c1als folder to begin with? It seems like we just weren't careful when organizing these screens.

Another problem one is the C1_FE_STATUS.adl screen.  Can I create a c1gds folder, and rename that screen to C1GDS_FE_STATUS.adl?  Objections?

 

  5792   Wed Nov 2 22:02:39 2011 JenneUpdatePhotosNew screen snapshot script written!

After lots of trial and error, and a little inspiration from Koji, I have written a new script that will run when you select "update snapshot" in the yellow ! button on any MEDM screen. 

Right now, it's only live for the OAF_OVERVIEW screen.  View snapshot and view prev snapshot also work. 

Next on the list is to make a script that will create the yellow buttons for each screen, so I don't have to type millions of things in by hand.

The script lives in:  /cvs/cds/rtcds/caltech/c1/scripts/MEDMsnapshots, and it's called....wait for it....... "updatesnap".

 

  5791   Wed Nov 2 21:49:59 2011 KatrinUpdateCDSc1iscey computer died again

while I was not doing anything on the machine.

  5790   Wed Nov 2 21:15:06 2011 KatrinUpdateCDSand again c1scy.mdl compiled

I changed an ADC channel for GCY_ERR and thus recompiled the c1scy model.

  5789   Wed Nov 2 20:56:49 2011 KatrinUpdateCDSdigital zeros at C1:X05-MADC0 (c1scy)

Channels C1:X05-MADC0_TP_XXX with XXX 2-9, 14-19, 21-27, 29-31 showed digital zeros.

Some of these channels are used in c1scy.mdl, e.g. for OSEM stuff. I guess this is not optimal...

  5788   Wed Nov 2 19:32:20 2011 kiwamuUpdateLSCPOP22/110 installed

[Steve / Kiwamu]

 The POP22/110 RFPD has been installed. It is PDA10A from Thorlabs instead of the usual home-made RFPD.
For an RF cable we rerouted one of the spare Heliax cables for it.
The cable is the one which used to be served for the 166MHz ASC wavefront sensors, picking up the RF source signal at 1X2 and sending it to the LSC rack.
 
 - - Remaining tasks - -
  + Fine alignment
  + Connection at the LSC rack
  + Update of the table diagram

Quote from #5783

They were traced and labeled. One goes to 1X2 and the other to AS-ISCT. They are Andrew Heliax 1/4" od. made by CommScone,  model number FSJ1-50A

  5787   Wed Nov 2 17:33:28 2011 KatrinUpdateGreen LockingADC channels for slow control

connector J9B of hardware ADC --> ch1 in software ADC --> GCY_ERR

connector J14 of hardware ADC --> ch11 in software ADC --> GCY_PZT

connector J15 of hardware ADC --> ch13 in software ADC --> GCY_REFL_DC

 

 

  5786   Wed Nov 2 17:29:10 2011 KatrinUpdateCDSc1scy.mdl compiled

Slight modification on that model:

  • terminated Q_out of Lockins to be able to compile the old model
  • assigned other ADC channels to GCY (green YARM)
  5785   Wed Nov 2 14:07:25 2011 ZachUpdateelogrestarted

Elog was hanging, restarted with the script.

  5784   Wed Nov 2 11:29:04 2011 not ZachUpdateSUS"Dr. SUS" progress

Quote:

Now that NDS2 is working, I was able to run all parts of diagAllSUS (except for the matrix writing) without errors. Barring issues with standard commands (freeswing, turnOnWatchDogs.csh, caput, etc.), Dr. SUS should be ready to go.

As requested, I have suppressed the generation of an elog post containing the results of the diagonalization.

Dr. SUS will run on allegra as a cron job at 8:00am on Sundays:

allegra:peakFit>crontab -l

PATH=/bin/:/usr/bin/:/cvs/cds/caltech/apps/linux64/matlab/bin/:/cvs/cds/caltech/apps/linux/gds/bin/:/cvs/cds/caltech/apps/linux/tds/bin/

PWD=/cvs/cds/caltech/apps/linux/gds/lib/

LD_LIBRARY_PATH=/cvs/cds/caltech/apps/linux/gds/lib

01 05 * * 0 cat /cvs/cds/caltech/users/kiwamu/work/20091026_OMC-LSC-diag/src/OMC_LSC_timing.m | matlab -nosplash -nodesktop > /dev/null

 

0 8 * * 0 /cvs/cds/rtcds/caltech/c1/scripts/SUS/peakFit/runDrSUS

Also: should these variable definitions really be here? 

Cron starts with only a very minimal PATH.  However, this shouldn't be an issue if you specify the full path to the commands.

The rest of the env vars should probably not be there.

Also, Zach:  using /cvs/cds is now forbidden.  We need to put this script in the right place.

  5783   Wed Nov 2 10:41:47 2011 steveUpdateGeneral2 spare heliax at the LSC rack

They were traced and labeled. One goes to 1X2 and the other to AS-ISCT. They are Andrew Heliax 1/4" od. made by CommScone,  model number FSJ1-50A

  5782   Wed Nov 2 10:04:05 2011 steveUpdateSUSMC2 chamber anchoring tightened

Quote:

 Thinks to do before the NEXT realignment:

B,  tie 4 ancher bolts on table legs to the floor

C,  tie 4 dog clamps between table and chamber

D,  check the locked position of the 4 x 4 positioning screws

E,  check bellow protecting 4 tubes are not shorting

A,  here is the concrete  slab cut

 

It reminds me to check the  IFO vacuum envelope dog clamps on the chambers to floor  with torque wrench.

 

 

 I locked the PMC and MC resonated in TM00 right on. Than I started checking anchoring screws with the torque wrench.

B,  3/8 foor bolts were tight, They were set  to 50 ft/lbs

C,  dog clamp bolts were tight at 80 ft/lbs

D,  horizontal position locks were lose. They were locked by finger.

E,  below covers were reset not to short

The MC is locking in TM01 mode now 

  5781   Wed Nov 2 01:53:08 2011 ZachUpdateSUS"Dr. SUS" progress

Now that NDS2 is working, I was able to run all parts of diagAllSUS (except for the matrix writing) without errors. Barring issues with standard commands (freeswing, turnOnWatchDogs.csh, caput, etc.), Dr. SUS should be ready to go.

As requested, I have suppressed the generation of an elog post containing the results of the diagonalization.

Dr. SUS will run on allegra as a cron job at 8:00am on Sundays:

 

 

allegra:peakFit>crontab -l

 

PATH=/bin/:/usr/bin/:/cvs/cds/caltech/apps/linux64/matlab/bin/:/cvs/cds/caltech/apps/linux/gds/bin/:/cvs/cds/caltech/apps/linux/tds/bin/

 

PWD=/cvs/cds/caltech/apps/linux/gds/lib/

 

LD_LIBRARY_PATH=/cvs/cds/caltech/apps/linux/gds/lib

 

01 05 * * 0 cat /cvs/cds/caltech/users/kiwamu/work/20091026_OMC-LSC-diag/src/OMC_LSC_timing.m | matlab -nosplash -nodesktop > /dev/null

0 8 * * 0 /cvs/cds/rtcds/caltech/c1/scripts/SUS/peakFit/runDrSUS

Also: should these variable definitions really be here?

 

 

  5780   Tue Nov 1 23:13:28 2011 ZachUpdateComputersNDS2 channel files

I did some messing around with the NDS2 config and channel files and things seem to be working as expected... for now. SENSOR channel data can be acquired for all sensors on all hanging optics.

What I did:

  • NDS2 gets its channel lists from .../users/jzweizig/nds2-mafalda/nds2.conf, which is called in the start-nds2 script. Within this, there are channel.file lines that specify which channels are available for raw and trend data. The four files that were listed were:
    • C-R-raw-channel_list.txt
    • C-M-ChanList.txt
    • C-T-ChanLIst.txt
    • C-R-online-channel_list.txt (this one was listed after a hashed line, which was suspicous---see below)
  • I noticed that a grep for SENSOR only returned lines for non-MC mirrors in both "R" files
  • I also noticed that calling NDS2_GetChannels('mafalda.martian:31200') did not return any non-suffixed (i.e., raw) channel names for MCx_SENSOR channels, while non-MC SENSOR channels each had two non-suffixed listings. I thought this was strange.
  • I manually added the line "C1:SUS-MC1_SENSOR_UL 0 real_4 2048 C-R" to one of the "R" channel files, then restarted the NDS2 server, and that channel was still not served. I figured that the second "R" channel file might have been left in the config file as a mistake, so I commented it out, restarted the NDS2 server, and was able to get MC1_SENSOR_UL data. I have left the comment-out there, with a signed EDIT.
  • Wary of (and too lazy to) manually add lines for all 5 sensors for each MC mirror, I decided to try generating a channel file using the most recent .gwf file in the frames, as indicated in Joe and John's elog post. To do this, while in .../nds2-mafalda/, I ran:
    • /cvs/cds/caltech/users/jzweizig/nds2-server/bin/buildChannelList /frames/full/10042/C-R-1004246528-16.gwf > C-R-raw-channel_list.txt
  • A grep for SENSOR in the new C-R-raw-channel_list.txt now returned lines for all MC mirror sensors... BUT NOT FOR ETMY(?!). I tried some slightly older .gwf files (all from today), but the ETMY files never showed up. I had no choice but to enter them manually. Another odd thing is that the channel file generated this way seems to be fairly jumbled up, in the sense that there is no clear top-town order (e.g. SUS-BS_blah then SUS-ETMX_blah). Instead, some SUS-BS channels are here, some are after SUS-ETMX or SUS-PRM channels, etc. Look at the file to know what I mean.
  • The original raw channel file is there as C-R-raw-channel_list.txt.bak.1004247481.

In any case, as I said, everything appears to be working now, but as soon as we try to generate a new channel file using the prescribed means, there will inevitably be channels omitted. Someone who knows more than me should get to the bottom of this and wiki a strict, detailed procedure for how this is to be done.

  5779   Tue Nov 1 19:26:38 2011 ZachUpdateSUS"Dr. SUS" progress

EDIT: Ran Jamie's NDS2 reset method and it's still not serving up MC data.

I am debugging the SUS input matrix diagonalization code (which I have affectionately termed "Dr. SUS") offline by running everything except the actual writing of the matrices (%writeSUSinmat()) in a sandbox directory. Some notes:

  • nodus doesn't know NDS2, but this isn't a problem given Kiwamu's reply, since we don't want to push results to the elog. I will still have the script push "optics kicked---don't touch" alerts to the elog, but I realized this can be done easily from a remote computer via in-script SSH commands. My next choice was pianosa, but even though she knows NDS2, she spits out MEX errors when I try to get data. Third stop was allegra, and she plays nice 
  • NDS2 is unable to read the MC mirror sensor data. Data from all other optics are retrieved without a problem. I verified that raw data was being written by plotting playback data in dataviewer, and I tried various times and the problem persisted. I suspect the problem is on the NDS2 server side:

The most recent entry in .../users/jzweizig/nds2-mafalda/channel_history shows nothing for MC1 2 or 3

 

mafalda:channel_history>more C-R-999986048-178368.cls | grep SENSOR
C1:SUS-BS_SENSOR_LL raw 0 real_4 2048 C-R
C1:SUS-BS_SENSOR_LR raw 0 real_4 2048 C-R
C1:SUS-BS_SENSOR_SIDE raw 0 real_4 2048 C-R
C1:SUS-BS_SENSOR_UL raw 0 real_4 2048 C-R
C1:SUS-BS_SENSOR_UR raw 0 real_4 2048 C-R
C1:SUS-ETMX_SENSOR_LL raw 0 real_4 2048 C-R
C1:SUS-ETMX_SENSOR_LR raw 0 real_4 2048 C-R
C1:SUS-ETMX_SENSOR_SIDE raw 0 real_4 2048 C-R
C1:SUS-ETMX_SENSOR_UL raw 0 real_4 2048 C-R
C1:SUS-ETMX_SENSOR_UR raw 0 real_4 2048 C-R
C1:SUS-ETMY_SENSOR_LL raw 0 real_4 2048 C-R
C1:SUS-ETMY_SENSOR_LR raw 0 real_4 2048 C-R
C1:SUS-ETMY_SENSOR_SIDE raw 0 real_4 2048 C-R
C1:SUS-ETMY_SENSOR_UL raw 0 real_4 2048 C-R
C1:SUS-ETMY_SENSOR_UR raw 0 real_4 2048 C-R
C1:SUS-ITMX_SENSOR_LL raw 0 real_4 2048 C-R
C1:SUS-ITMX_SENSOR_LR raw 0 real_4 2048 C-R
C1:SUS-ITMX_SENSOR_SIDE raw 0 real_4 2048 C-R
C1:SUS-ITMX_SENSOR_UL raw 0 real_4 2048 C-R
C1:SUS-ITMX_SENSOR_UR raw 0 real_4 2048 C-R
C1:SUS-ITMY_SENSOR_LL raw 0 real_4 2048 C-R
C1:SUS-ITMY_SENSOR_LR raw 0 real_4 2048 C-R
C1:SUS-ITMY_SENSOR_SIDE raw 0 real_4 2048 C-R
C1:SUS-ITMY_SENSOR_UL raw 0 real_4 2048 C-R
C1:SUS-ITMY_SENSOR_UR raw 0 real_4 2048 C-R
C1:SUS-PRM_SENSOR_LL raw 0 real_4 2048 C-R
C1:SUS-PRM_SENSOR_LR raw 0 real_4 2048 C-R
C1:SUS-PRM_SENSOR_SIDE raw 0 real_4 2048 C-R
C1:SUS-PRM_SENSOR_UL raw 0 real_4 2048 C-R
C1:SUS-PRM_SENSOR_UR raw 0 real_4 2048 C-R
C1:SUS-SRM_SENSOR_LL raw 0 real_4 2048 C-R
C1:SUS-SRM_SENSOR_LR raw 0 real_4 2048 C-R
C1:SUS-SRM_SENSOR_SIDE raw 0 real_4 2048 C-R
C1:SUS-SRM_SENSOR_UL raw 0 real_4 2048 C-R
C1:SUS-SRM_SENSOR_UR raw 0 real_4 2048 C-R

 

  5778   Tue Nov 1 18:45:48 2011 JenneUpdateComputersAllegra's screens

I was trying to give Allegra a second head, and it didn't quite work.  It's still in progress.  Steve, you might not like how I've 'mounted' the second monitor, so we can talk about something that might work tomorrow.

  5777   Tue Nov 1 18:16:50 2011 DenUpdateAdaptive FilteringAdaptive filter witness and EP SNR

Quote:

 

Coherence of seismometers to MCL:


STS1 is located at the vertex. x-axis along the x arm.
GUR1 is located at the IMC MC2 mirror. Same orientation.
Coherence.png

=> 1. Only the x-direction has good coherence (to be expected)
     2. Only good coherence at 1.5-4Hz (huh?)

So probably other noise sources are dominating. Let's look into noise projections. Remember IMC autoalignment is off.

A quick adaptive filter run with only the GUR1 and STS1 witnesses applied only to MCL didn't really do anything. Some more thought needs to be invested into the AA and shaping filters.

The possible explanation to this effect is the following:

Seismic noise mainly consists of the Love and Rayleigh surface waves. In the simulations we've taken 2 perpendicular Love waves and 2 perpendicular Rayleigh waves that interfere under the test mirrors. This interference produces both translational and tilt movements. Then we can see the coherence between translational motion and cavity length.

translation_length.jpg

1. The coherence at big frequencies is small due to the passive isolation.

2. The coherence at 1 Hz is 0 due to the wire resonance.

3. The coherence between 1 and 10 Hz is reasonable. At the real 40m's measurements we can see only good coherence for gur1_x and sts1_x but this is the matter of adjusting seismic waves amplitude and direction. In the simulation we've assumed that all waves are of the same amplitude. The really interesting thing is that

4. The coherence below 0.8 Hz began to grow. We don't see this in real measurements.

But let's simulate the seismometer measurements. It measures not only translational motion but also tilt and with amplitude proportional to g / omega^2. On the Figure below the spectrum of translation motion, tilt and tilt as seen by seismometer is presented. We can see that at low frequencies tilt begins to dominate over the translational motion. We assumed the speed of waves in the region 30 - 60 m/sec.

trans_tilt.jpg

Let's now plot the coherence between the cavity length and seismometer signal.

seismic_length.jpg

We can see that the coherence between seismic signal from measured by seismometer and cavity length is gone below 1 Hz where tilt becomes important.

Now let's try to filter out the seismic noise from the cavity length using both static Wiener filtering and adaptive Mfxlms algorithm. For both filters we've used AA filter before the filters and also AI filter after adaptive filter. The downsampling ratio was 4, the sample frequency 256. We can see that nothing is really subtracted due to the pollution of the seismometer signal due to tilt motion.

tilt_filtering.jpg

Assume we do the same computational experiment but with the seismometers that measure only ground translational motion and tilt do not affect on them. Then we have a reasonable subraction of seismic noise at low frequencies even with the filters of the length 100 as shown on the figure below.

filtering.jpg

Let's look through an order of magnitude analysis. Assume ground motion consists of only one wave with amplitude A and only vertical movement:  z(t) = A*sin(2 pi 0.1 t). So the frequency of the wave is 0.1 Hz. If A = 10-6 m => the amplitude of the suspended mirror motion is also approximately 10-6 m, as we have no isolation at low frequencies. The tilt angle has the amplitude alpha = 2*pi*A/lambda, where lambda - wavelength of the ground wave, lambda = v/f = 40/0.1 = 400 m, v - speed of the wave, f - frequency. Then alpha = 10-8 rad. If the distance between ground and mirror suspension point is 1 m, then mirror motion amplitude due to tilt is B = 10-8 m << A. 
It turns out that tilt does not effect much on the cavity length compared to the ground translational motion, but it affects a lot on the seismometer signals, that are used as witness signals in the filtering. For that reason we need tiltmeters to filter seismometer signals in order to obtain pure translational ground motion.
  5776   Tue Nov 1 16:02:15 2011 kiwamuUpdateSUSRe: SUS input matrix diagonalizer REMOVED from crontab

Quote:

It turns out that nodus doesn't know how to NDS2, so I can't run diagAllSUS as a cron job on nodus. To further complicate things, no other machines can run the elog utility, so I am going to have to do something shifty...

 Actually in the last 40m meeting we discussed the SUS diagnosis and decided not to post the results on the elog.

Alternatively we will have a summary web page which will contain all the information (sensitivity, UGF monitors, etc.,) and will be updated everyday like GEO.

This will be a place where we should post the SUS results every week.

So we don't need to worry about the cron-elog job, and for running the scripts you can simply use one of the lab machines as a cron host.

Once you get the scripts running on a machine as a cron job, it will be the point where we should quit developing the SUS diagonalizer and pass it to the web summary people.

  5775   Tue Nov 1 13:46:03 2011 ZachUpdateSUSSUS input matrix diagonalizer REMOVED from crontab

It turns out that nodus doesn't know how to NDS2, so I can't run diagAllSUS as a cron job on nodus. To further complicate things, no other machines can run the elog utility, so I am going to have to do something shifty...

  5774   Tue Nov 1 13:41:38 2011 MirkoUpdateLSCAM modulation due to non-optimal SB frequency

Quote:

[Kiwamu, Mirko]

Non-optimal 11MHz SB frequency causes PM to be transformed into AM.
m_AM / m_PM = 4039 * 1kHz / df , with df beeing the amount the SB freq. is off.

Someone might want to double check ths.

 Actually there was an error.

For 11MHz it is:
m_AM / m_PM = 2228 * 1kHz / df

For 55MHz:
m_AM / m_PM = 99.80 * 1kHz / df

see PDF

Attachment 1: IMC.pdf
IMC.pdf
  5773   Mon Oct 31 21:46:32 2011 kiwamuUpdateLSCDependence of Recycling gain on incident beam pointing

I horizontally swept the translation of the incident beam in order to investigate a possible clipping in Power-recycled Michelson (PRMI).

The recycling gain of PRMI depended on the beam pointing but it did't improve the recycling gain.

I guess the amount of the entire shift I introduced was about +/- the beam diameter = +/- 5 mm or so.

Within the range of about +/- 5mm in the horizontal beam translation I didn't find any obvious sign of a clipping.

 

(Measurement)

 This is the procedure which I did:
  (1) Some amount of offsets were introduced on MC2 in both PIT and YAW such that the PZT1 won't rail (#5762).
      => Every time when I introduced the offset I realigned the zig-zag mirrors on the PSL table to maintain the high transmissivity of MC.
  (2) Fine tuning of the MC offsets so that the PZT1_X EPICS value becomes almost zero when the beam is aligned down to the Y arm.
     => 0.523 in C1:ASC_PZT1_X became a point where the coupling of the beam into the Y arm was maximized.
     => Last time the direction which we investigated was the positive side from this zero point (#5709) in PZT1_X.
  (3) Aligned MICH by steering BS.
  (4) Locked PRMI with carrier resonating and aligned PRM to maximized the power recycling gain which was obtained from POYDC.
  (5) Translated the beam pointing
     => First I shifted PZT1_X by a wanted amount.
    => Then I locked the Y arm and realigned PZT2_X by maximizing the Y arm transmission.
          This procedure should give us a pure translation on the incident beam.
  (6) Repeated the same procedure (3) through (5) in each PZT1 position.
 

(Results)

Here shows the measured recycling gain and the power reflectivity of PRMI as a function of the beam pointing.

beam_scan_PRCL.png

 Upper plot : measured recycling gains (Red) observed maximum values (Black) measured values on average.
 Lower plot : measured power reflectivity of PRMI (Blue) observed minimum values (Black) measured values on average.
 
 As shown in the plots the recycling gain could go up to 8 at some points.
As the PZT went away from 0 it decreased and eventually became about 3 in each side.
The reflectivity showed the minimum value of 0.4 when the PZT1 was at -1 in EPICS value.
One hypothesis to explain this plot can be that : we are just seeing the effect of the incident beam misalignment.

 

  5772   Mon Oct 31 19:39:00 2011 JenneUpdateAdaptive FilteringScreens, code, computers

[Mirko, Jenne]

I finished (mostly? maybe?) the OAF screens.  They're pretty awesome. 

While we were playing with the OAF, trying to do some oafing, the output of the code decided to just be zeros.  We did a test, and in the c-code set the output to be a constant value, and that worked.  But when we put it back to normal (being adaptive) and recompiled, it still only outputs zeros.  The code is receiving both witness and error signals, so we don't know what's going on.

In the course of trying to make things work again, we did a complete reboot of c1lsc and c1sus.  Did a burt restore.  Everything (except our weird code problem) should be fine again.  Optics are damping happily.

  5771   Mon Oct 31 15:58:25 2011 steveUpdateSUSthinking of black glass baffles for 40m

Shade 12 black glass-green lines: 1/8" thick,  6" wide  x 7 " high  mounted on angle bracket below.  Aperture 1" diameter. Let me know what do you think.

 

Attachment 1: 03260901.PDF
03260901.PDF
Attachment 2: 03260902.PDF
03260902.PDF
  5770   Mon Oct 31 14:06:16 2011 ZachUpdateSUSSUS input matrix diagonalizer added to crontab

I actually tried to do this last night, but I was getting a terminal error from nodus. Jamie told me to just manually redefine the TERM variable and it would work. So, I have added kickAll and diagAllSUS to controls@nodus's crontab:

 

nodus:~>crontab -l

0 5 * * * /opt/rtcds/caltech/c1/scripts/backup/rsync.backup

0 7 * * * /opt/rtcds/caltech/c1/scripts/backup/check_backup.sh 

0 17 * * 0 /cvs/cds/caltech/users/jenne/LIGOX/LIGOX_Pizza_Reminders.sh

0 8 * * 0 /cvs/cds/rtcds/caltech/c1/scripts/SUS/peakFit/kickAll

0 18 * * 0 /cvs/cds/rtcds/caltech/c1/scripts/SUS/peakFit/diagAllSUS

Again, their functionality should be checked before this is allowed to run on Sunday!

 

ELOG V3.1.3-